Find all the solutions of 6cos x - 3√3 = 0
1
Expert's answer
2012-05-29T10:27:05-0400
Find all the solutions of 6cos x - 3√3 = 0.

6cos x - 3√3 = 0 ==>
6cosx = 3√3 ==>
cosx = √3/2 ==>
cosx = π/6 (= 30°) and 11π/6 ( = 330°).